The demand and supply curves for a product are given in terms of price, p, by q = 2600 - 20p and q = 10p - 400

A. Find the equilibrium price and quantity.

B. A specific tax of $12 per unit is imposed on suppliers. Find the new equilibrium price and quantity. The new equilibrium price (including tax) is $______ and the new equilibrium quantity is ______ units.

C. How much of the $12 tax is paid by consumers and how much by producers? The tax paid by the consumer is $______ and the tax paid by the producer is $_____ of the tax.

D. What is the total tax revenue received by the government? $________
Answer #1

A) Equilibrium condition; Qd = Qs

2600 - 20p = 10p - 400

3000 = 30p

p = 100

q = 2600 - 20(100) = 2600 - 2000

q = 600

B) When tax is imposed on seller; Qs = 10(p - 12) - 400

Qs' = 10p - 120 - 400

Qs' = 10p - 520

New Equilibrium; Qd = Qs'

2600 - 20p = 10p - 520

3120 = 30p

p = 104

q = 10(104) - 520 = 1040 - 520

q = 520

C) Tax paid by consumer = 104 - 100 = $ 4

Tax paid by seller = 100 - 92 = $ 8

D) Tax revenue = (104 - 92) x 520 = 12 x 520 = 6240

ax Tax on ๑ 600